CS2(g) + O2(g) CO2(g) + SO2(g)

Consider the unbalanced equation above. A 0.150 g sample of CS2 is reacted with 285 mL of O2 at STP. What is the total volume of the final mixture of gases at STP?
1
Expert's answer
2019-09-27T04:46:32-0400

Given equation is "CS_2+O_2\to CO_2+SO_2"

Balanced equation will be "CS_2+3O_2\to CO_2+2SO_2"

As 1 mole of "CS_2" reacts with 3 moles of "O_2" it gives 1 mole of "CO_2" and 2 moles of "SO_2"

Molar mass of "CS_2" "=" 76"g\/mol"

Number of moles of "CS_2 =\frac{150}{76}=0.001974"

Number of moles of "O_2" "=\frac{285}{22400}=0.0127"

Number of moles of "O_2" required to completely react with "CS_2" "=3\times0.001974=0.005922"

Number of unreacted moles of "O_2=0.0127-0.005922=0.006778"

Number of moles of "CO_2" formed"=0.001974"

Number of moles of "SO_2" formed"=2\times0.001974=0.003948"

After reaction total number of gaseous moles"=0.006778+0.001974+0.003948=0.0127"

Total volume of final mixture "=0.0127\times22400=285ml"